# thanks: Corresponding author

### v1

@@ Abstract
We study the non-leptonic decays of $\Omega_{b}\rightarrow\Omega_{c}P(V)$ with eight pseudoscalar and vector mesons using the QCD factorization approach. We analyze all relevant topologies (the tree-level, color-suppressed, and penguin) of these decays and calculate the decay amplitude for each separately. We determine the decay rates, branching ratios and compare our results with those from other theoretical predictions. The results obtained may be useful for the analysis of the related data in both ongoing and future experiments.
@@ I Introduction
In Fig.2, the penguin diagrams relevant to these decays are shown; they can hadronize into $D^{(*)}$ and $D_{s}^{(*)}$ mesons. Consequently, their occurrence probability is lower than that of tree-level topologies, and they do not significantly modify the total decay rate. Nevertheless, they can generate additional complex phases that participate in amplitude interference and therefore play a crucial role in the size of CP violation. Theoretical analyses of non-leptonic decays rely on factorization theorems, which constitute fundamental tools within perturbative QCD frameworks. These theorems separate physics across different energy scales, distinguishing perturbatively computable short-distance interactions from non-perturbative long-distance effects. The factorization theorem extends the Euclidean operator product expansion to time-like kinematics, with rigorous proofs derived from perturbative QCD through Feynman diagram analysis in the Collins-Soper-Sterman formalism Catani:1989ne ; Collins:1981uk ; Collins:1984kg . Heavy quark masses further suppress strong interaction effects through systematic expansions in heavy quark effective theory. Naive factorization was initially applied to B-meson decays and later generalized to heavy-baryon systems. While naive factorization performs well for color-allowed, tree-dominated processes, it fails to adequately describe color-suppressed or penguin-dominated channels. In such cases, nonfactorizable QCD corrections, arising from interactions between the emitted hadron and the recoiling system, become essential. The general factorization formula for separating the hadronic matrix element is given as follows Lu:2009cm ; Beneke:2006thn :
$\langle h_{1}h_{2}|{\cal O}_{i}|B\rangle=\Phi_{h_{2}}(u)\otimes\left(T^{I}(u)F^{Bh_{1}}(0)+C^{II}(\tau,u)\otimes\Xi^{Bh_{1}}(\tau,0)\right).$ (2)
Factorization involves the QCD form factor $F^{Bh_{1}}(0)$ alongside an unknown non-local form factor $\Xi^{Bh_{1}}(\tau,0)$, $\Phi_{h_{2}}$ is light-cone distribution amplitudes. Different parts of the factorization formula for non-leptonic two-body decays have been treated in several ways, leading to three major theoretical approaches: perturbative QCD (pQCD), QCD factorization (QCDF), and Soft-Collinear Effective Theory (SCET). The PQCD approach is based on transverse-momentum-dependent factorization and both form factors $F^{Bh_{1}}(0)$ and the non-local form factor $\Xi^{Bh_{1}}(\tau,0)$ are assumed to be dominated by short-distance dynamics and are therefore calculable using, $k_{\perp}$-factorization. All four quantities in the factorization formula, including $T_{I}$ and $C_{II}$, have been calculated at leading order, with some next-to-leading order corrections recently included for $T_{I}$.
The QCDF approach, founded on collinear factorization, has been extensively applied to a wide range of two-body nonleptonic B-meson decay channels, and most of its predictions for decay rates are consistent with experimental measurements. The QCD factorization framework can be more elegantly reformulated in the language of SCET. It works reliably at leading power in the heavy-quark expansion and can calculate non-factorizable QCD corrections Neubert:2000ag . In this approach, the standard heavy-to-light form factors $F^{Bh_{1}}(0)$ are assumed to receive significant soft (long-distance) contributions, rendering them incalculable in perturbation theory. In contrast, $\Xi_{Bh_{1}}(\tau,0)$ is dominated by perturbative hard-collinear interactions and factorizes further into light-cone distribution amplitudes. Within QCDF, the form factor $F^{Bh_{1}}(0)$ is treated as a phenomenological input, typically taken from QCD sum rules or lattice QCD calculations. In this study, we employ the QCD factorization approach, which enables a comprehensive calculation of all topologies relevant to these decays. Form factors have been calculated using the QCD Sum Rules (QCDSR) method, which is one of the powerful and predictive models in the non-perturbative regime of QCD. This method has provided good predictions and results comparable to experiments so far, making it a reliable non-perturbative approach for studying hadronic decays.
@@ II The Decay Amplitude
where $M$ denotes the meson (pseudoscalar or vector). In the next step, using the factorization method, the hadronic matrix element is separated into two distinct parts:
$\l \left\langle\Omega_{c}M\left|O_{i}\right|\Omega_{b}\right\rangle\,\to\langle{M}|\mathcal{J}^{\mu}_{ud}|0\rangle\otimes\langle{\Omega_{c}}|\mathcal{J}_{\mu,cb}|{\Omega_{b}}\rangle\,$, (6)
@@ III DECAY RATES
$\Omega_{b}\rightarrow\Omega_{c}D_{s}$ T,C,P $V_{bc}V_{cs}(a_{1}+a_{2})-V_{tb}V_{td}(a_{4}+a_{10}+R(a_{6}+a_{8}))]$
$\Omega_{b}\rightarrow\Omega_{c}D_{s}^{*}$ T,C,P $V_{bc}V_{cs}(a_{1}+a_{2})-V_{tb}V_{td}(a_{4}+a_{10})]$
@@ IV Numerical Analysis
By substituting the form factors into the decay rate expressions given in III, the decay rates and branching ratios for the pseudoscalar($\pi,K,D,D_{s}$) and vector ($\rho,K^{*},D^{*},D_{s}^{*}$) mesons are obtained and reported in Tables 5 and 6. These tables include the other predictions in the literature, as well. The obtained decay rates and branching ratios, including the uncertainties, are overall in good agreement with the other existing predictions. It is worth mentioning that different references consider different topological contributions: The differences among the presented results can be attributed to this fact. References Li:2021kfb ; Zhao:2018zcb and Chua:2019yqh ; Chua:2019lgb calculate the related form factors using the light-front method and examine nonleptonic decays only at the tree level (color allowed). References Gutsche:2018utw ; Ivanov:1997hi employ the covariant confined quark model and investigate the decays at the tree level (color allowed), incorporating QCD penguin operators. Reference Cheng:1996cs uses the non-relativistic quark model to determine form factors and calculates the decay rates at the tree level including color-suppressed contributions. The Reference Patel:2025gbw employs the independent quark model within a relativistic framework and examines the corresponding decays at the tree level (color allowed). In contrast, the present work provides the most complete analysis of the nonleptonic decay $\Omega_{b}\to\Omega_{c}$ incorporating tree-level, color-suppressed, and penguin topologies. To illustrate the impact of each topology on the decay rates, the partial decay rates corresponding to individual topological contributions for each decay channel are listed in Table 7. This decomposition allows for a clear assessment of the relative importance of different topologies across various decay modes. As shown, although the pure decay rates from color-suppressed and penguin diagrams are small, they can significantly affect the total decay rates and should not be neglected. These results provide a comprehensive reference for analyzing asymmetry effects and their role in different decay channels from the viewpoint of CP violation effects.
@@ Appendix A The penguin-operator coefficients
$C^{AB}=\sum_{a}\mathrm{Tr}(\Gamma^{A}_{a}\Gamma^{B}_{b}\Gamma^{A}_{a}\Gamma^{B}_{b}).$ (32)
After a Fierz transformation, it takes the form below:

### v2

@@ Abstract
We study the non-leptonic decays of $\Omega_{b}\rightarrow\Omega_{c}P(V)$ with eight pseudoscalar and vector mesons using the naive factorization approach. We analyze all relevant topologies (the tree-level, color-suppressed, and penguin) of these decays and calculate the decay amplitude for each separately. We determine the decay rates, branching ratios and compare our results with those from other theoretical predictions. The results obtained may be useful for the analysis of the related data in both ongoing and future experiments.
@@ I Introduction
In Fig.2, the penguin diagrams relevant to these decays are shown; they can hadronize into $D^{(*)}$ and $D_{s}^{(*)}$ mesons. Consequently, their occurrence probability is lower than that of tree-level topologies, and they do not significantly modify the total decay rate. Nevertheless, they can generate additional complex phases that participate in amplitude interference and therefore play a crucial role in the size of CP violation. Theoretical analyses of non-leptonic decays rely on factorization theorems, which constitute fundamental tools within perturbative QCD frameworks. These theorems separate physics across different energy scales, distinguishing perturbatively computable short-distance interactions from non-perturbative long-distance effects. The factorization theorem extends the Euclidean operator product expansion to time-like kinematics, with rigorous proofs derived from perturbative QCD through Feynman diagram analysis in the Collins-Soper-Sterman formalism Catani:1989ne ; Collins:1981uk ; Collins:1984kg . Heavy quark masses further suppress strong interaction effects through systematic expansions in heavy quark effective theory. Naive factorization was initially applied to B-meson decays and later generalized to heavy-baryon systems. Naive factorization assumes that the hadronic matrix element of a four-quark operator can be written as the product of two single-hadron matrix elements, neglecting nonfactorizable effects such as hard-gluon exchange and spectator interactions.
In this study, we employ the naive factorization approach, which enables a comprehensive calculation of all topologies relevant to these decays. Form factors have been calculated using the QCD Sum Rules (QCDSR) method, which is one of the powerful and predictive models in the non-perturbative regime of QCD. This method has provided good predictions and results comparable to experiments so far, making it a reliable non-perturbative approach for studying hadronic decays.
@@ II The Decay Amplitude
where $M$ denotes the meson (pseudoscalar or vector). In the next step, using the naive factorization method, the hadronic matrix element is separated into two distinct parts:
$\left\langle\Omega_{c}M\left|O_{i}\right|\Omega_{b}\right\rangle\,\to\langle{M}|\mathcal{J}^{\mu}_{ud}|0\rangle\otimes\langle{\Omega_{c}}|\mathcal{J}_{\mu,cb}|{\Omega_{b}}\rangle\,$, (5)
@@ III DECAY RATES
[…]
@@ IV Numerical Analysis
By substituting the form factors into the decay rate expressions given in III, the decay rates and branching ratios for the pseudoscalar($\pi,K,D,D_{s}$) and vector ($\rho,K^{*},D^{*},D_{s}^{*}$) mesons are obtained and reported in Tables 5 and 6. These tables include the other predictions in the literature, as well. The obtained decay rates and branching ratios in the present study, considering the uncertainties, are overall in good agreement with the other existing predictions. It is worth mentioning that different references use different form factors calculated via different approaches and consider different topological contributions: The differences among the presented results for some channels can be attributed in the first stage to the different form factors used in these references as well as different topologies considered. The choice of form factors and their behavior in terms of $q^{2}$ as well as their values at $q^{2}=m^{2}_{meson}$ play important role in enhancing our results for the branching ratios in some channels in Table 6. References Li:2021kfb ; Zhao:2018zcb and Chua:2019yqh ; Chua:2019lgb calculate the related form factors using the light-front method and examine nonleptonic decays only at the tree level (color allowed). References Gutsche:2018utw ; Ivanov:1997hi employ the covariant confined quark model and investigate the decays at the tree level (color allowed), incorporating QCD penguin operators. Reference Cheng:1996cs uses the non-relativistic quark model to determine form factors and calculates the decay rates at the tree level including color-suppressed contributions. The Reference Patel:2025gbw employs the independent quark model within a relativistic framework and examines the corresponding decays at the tree level (color allowed). In contrast, the present work provides the most complete analysis of the nonleptonic decay $\Omega_{b}\to\Omega_{c}$ incorporating tree-level, color-suppressed, and penguin topologies. To illustrate the impact of each topology on the decay rates, the partial decay rates corresponding to individual topological contributions for each decay channel are listed in Table 7. This decomposition allows for a clear assessment of the relative importance of different topologies across various decay modes. As shown, although the pure decay rates from color-suppressed and penguin diagrams are small, they can significantly affect the total decay rates and should not be neglected. These results provide a comprehensive reference for analyzing asymmetry effects and their role in different decay channels from the viewpoint of CP violation effects.
@@ Appendix A The penguin-operator coefficients
$C^{AB}=\frac{1}{16N_{B}}\mathrm{Tr}(\Gamma^{A}_{a}\Gamma^{B}_{b}\Gamma^{A}_{a}\Gamma^{B}_{b}),$ (31)
where $N_{B}$ is the number of gamma matrices in each group $B$: In our case $N_{B}=4$ . After a Fierz transformation, it takes the form below: